The father’s age is four times that of his son. Before 8 years, the father’s age was 16 times that of the son . Find the father’s age.

Answer by sachi(548)

he father’s age is four times that of his son.
let son is x years
so father is 4x yrs
Before 8 years,
son was x-8 years
& father was 4x-8 yrs
the father’s age was 16 times that of the son .
so 4x-8=16(x-8)=16x-128
or 16x-4x=128-8=120
or 12x=120
or x=10
so the father’s age=4*10=40 yrs
ans